# Enabling stratified sampling in high dimensions via nonlinear dimensionality reduction

### v1

@@ 5 Conclusion
The idea presented in this work could be extended in multiple directions. First, we applied NeurAM-based stratification to multifidelity Monte Carlo estimators, but other variance reduction strategies could also be considered. In addition, similarly to what we do in Section 4.3 for active subspaces, alternative techniques for dimensionality reduction could be leveraged to improve stratified sampling. Finally, this study can be extended to higher-order statistical moments of a given quantity of interest, or for variance reduction in sensitivity analysis.
@@ Acknowledgements
This article has been authored by an employee of National Technology & Engineering Solutions of Sandia, LLC under Contract No. DE-NA0003525 with the U.S. Department of Energy (DOE). The employee owns all right, title and interest in and to the article and is solely responsible for its contents. The United States Government retains and the publisher, by accepting the article for publication, acknowledges that the United States Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this article or allow others to do so, for United States Government purposes. The DOE will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan https://www.energy.gov/downloads/doe-public-access-plan. DES acknowledges support from NSF CDS&E award #2104831, and NSF CAREER award #1942662. AZ is supported by “Centro di Ricerca Matematica Ennio De Giorgi” and the “Emma e Giovanni Sansone” Foundation, and is member of INdAM-GNCS.

### v2

@@ 5 Conclusion
The idea presented in this work could be extended in multiple directions. First, we applied NeurAM-based stratification to multifidelity Monte Carlo estimators, but other variance reduction strategies could also be considered. In addition, similarly to what we do in Section˜4.3 for active subspaces, alternative techniques for dimensionality reduction could be leveraged to improve stratified sampling. One limitation of the current approach, inherited from the NeurAM methodology [ZGS25], concerns its performance for models exhibiting discontinuities and bifurcations in which the low-dimensional manifold is not simply connected, and addressing this limitation is an interesting direction for future work. Moreover, even if, ideally, a one-dimensional manifold should be sufficient to capture the response surface of a scalar quantity of interest, for more complex problems, a NeurAM with more than one dimension could provide improved results. However, this direction involves several nontrivial challenges. First, by moving to higher dimensions, we lose the interpretation that the strata tend to follow the contour lines of the model, and it becomes unclear how to perform stratification in the latent space so that it remains effective in the original domain. Second, inverse transform sampling is no longer straightforward in dimensions greater than one. Hence, one would need a method to map the latent space distribution to a uniform distribution on the unit hypercube, e.g., via normalizing flows, flow matching, or diffusion models, which should be trained in addition to NeurAM. Third, the extension of NeurAM itself to a multi-dimensional latent space requires investigation. In particular, as discussed in [ZGS25, Section 5], we would like to preserve some ordering or structure in the latent coordinates, similar to the active subspace method, where reduced variables are ranked based on their importance. Finally, this study can be extended to higher-order statistical moments of a given quantity of interest, or for variance reduction in sensitivity analysis.
@@ Acknowledgements
We thank the Editor and the anonymous referees for their thorough reading and for their insightful comments and valuable suggestions, which have helped improve and clarify this manuscript. This article has been authored by an employee of National Technology & Engineering Solutions of Sandia, LLC under Contract No. DE-NA0003525 with the U.S. Department of Energy (DOE). The employee owns all right, title and interest in and to the article and is solely responsible for its contents. The United States Government retains and the publisher, by accepting the article for publication, acknowledges that the United States Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this article or allow others to do so, for United States Government purposes. The DOE will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan https://www.energy.gov/downloads/doe-public-access-plan. DES acknowledges support from NSF CDS&E award #2104831, and NSF CAREER award #1942662. AZ is supported by “Centro di Ricerca Matematica Ennio De Giorgi” and the “Emma e Giovanni Sansone” Foundation, and is member of INdAM-GNCS. The authors would like to thank the Center for Research Computing at the University of Notre Dame for providing computational resources and support that were essential to generate the results for the Darcy flow problem.